# FREA: Feasibility-Guided Generation of Safety-Critical Scenarios with Reasonable Adversariality

### v1

@@ 4 Experiments
Following the experimental setup (Section 4.1), we conducted several experiments to affirm: The LFR of the AV is reasonable through offline training (Section 4.2); FREA effectively generates near-miss events while maintaining AV’s feasibility (Section 4.3); FREA shows generalization in AV testing across various surrogate AVs and environments (Section 4.4); and FREA continuously creates adversarial yet AV-feasible scenarios in complex traffic (Section 4.5).
@@ 4.1 Experiment Setup
Environment. We developed our environment using the Carla Simulator [24], as per SafeBench standards [25]. We combined “Scenario6” and “Scenario7” in SafeBench into a new “Scenario9” spread across “Town05” and “Town02”. In “Scenario9”, the AV must navigate predefined routes through different types of junctions while avoiding randomly placed background vehicles (BVs). To implement reasonable adversarial attacks, we employed the FREA method on selected CBVs, aiming to attack the surrogate AV with a goal-based adversarial reward, thus generating adversarial yet AV-feasible, safety-critical scenarios. More experimental details are provided in Appendix B.
@@ 4.2 Largest Feasible Region Learning of AV
Figure 2: Visualization of well-trained LFR $\left\{s^{\text{AV}}|V^{\ast}_{h}(s^{\text{AV}})\leq 0\right\}$ in various traffic scenarios. Red Rectangle: AV. Blue Rectangle: BV
According to the loss functions for the feasible value functions detailed in Eqs. 3 and 4, learning incorporates the constraint violation function $h(s^{\text{AV}})$. A straightforward definition of $h(s^{\text{AV}})$ is $\min_{i\in\{1,\ldots,N\}}d(\text{AV},\text{BV}^{i})$, where $d(\text{AV},\text{BV}^{i})$ represents the minimum distance between the bounding boxes of AV and ${\text{BV}}^{i}$. However, since the minimum distance between boxes is non-negative, we adopt a sparse design for $h(s^{\text{AV}})$, which helps to maintain a stable training balance of positive and negative samples, as advised by [15]:
$h(s^{\text{AV}}):=\begin{cases}-1&\text{if }\min_{i\in\{1,\ldots,N\}}d(\text{AV},\text{BV}^{i})>d_{th}\\ M&\text{if }\min_{i\in\{1,\ldots,N\}}d(\text{AV},\text{BV}^{i})\leq d_{th}\end{cases},$ (9)
@@
Figure 2 depicts the well-trained LFR of the AV in varying traffic scenarios. Figure 2(a) and (b) demonstrate that the infeasible region expands as the AV’s speed increases, indicating higher collision risks. Figure 2(c) and (b) confirm the LFR’s consistency across different scenarios. These cases show that the pre-trained LFR performs reliably under various traffic conditions, providing a strong foundation for CBV policy training.
@@ 4.3 Quantitative Analysis of Generated Scenarios
Evaluation of Near-miss Events. Research indicates that all safety-related incidents share common origins with near-misses [27, 28, 29], underscoring the importance of analyzing near-miss events in the generation of safety-critical scenarios. In line with [4], we use time-to-collision (TTC) and post-encroachment time (PET) as metrics to quantify these near-miss events.
@@
Figure 3(a) and (b) show the distributions of TTC and PET across different CBV methods. The PPO produces the most adversarial scenarios across all maps. Conversely, FPPO-RS and FREA present similar near-miss metrics but with a higher frequency of near-miss than Standard traffic flow. Given our focus on AV-feasible near-miss events, further analysis of AV’s feasibility is required.
Evaluation of AV’s Feasibility. While near-miss events are crucial for evaluating safety-critical scenarios, it’s essential to prevent over-collision events caused by excessive adversarial CBV policy. To ensure that FREA effectively minimizes over-collision events, we analyze the collision trajectories under different CBV methods with four key metrics: (1) Collision Rate (CR); (2) Infeasible Ratio (IR); (3) Infeasible Distance (ID); and (4) Collision Severity, including collision speed and impulse. (see Appendix D for details)
@@
As shown in Table 1, without the feasibility constraint, the PPO method leads to many severe collisions with high infeasible rates and longer infeasible distances from AV, indicating excessive adversarial behavior, as also seen in Figure 4. Conversely, under the feasibility constraint, both FPPO-RS and FREA mitigate the severity of these events. However, as for AV’s feasibility, FPPO-RS struggles with hard constraints, causing many infeasible events. In contrast, FREA strikes a better balance between adversarial and AV-feasible, resulting in the least severe collisions.
@@
To evaluate the generalization of FREA across various surrogate AVs and environments in AV testing, we assessed the driving performance of PlanT [30] and Expert [26] under different CBVs that were pre-trained with specific surrogate AVs on “Town05” and “Town02”. The performance metrics follow the setting in [25], where the OS represents the overall score (details in Appendix E).
Table 2 illustrates that the PPO method, pre-trained with various surrogate AV methods, exhibits notable variance in AV testing. This variance suggests that, without feasibility constraints, the CBV’s adversarial policy closely depends on the specific surrogate AV method. In contrast, FREA demonstrates remarkable stability in AV testing. This stability is primarily due to goal-based adversarial rewards and feasibility constraints, both independent of the specific surrogate AV method employed. Furthermore, FREA consistently aligns its relative AV testing results with standard traffic flows, suggesting that FREA facilitates critical events while ensuring unbiased relative AV testing results.
@@ 5 Conclusion
In conclusion, FREA employs the concept of the Largest Feasible Region (LFR) of the AV as the criterion for defining the upper bound for adversariality in safety-critical scenario generation. This approach aims to maximize adversariality while ensuring AV’s feasibility. Specifically, FREA employs a two-stage framework that initially pre-trains the AV’s optimal feasible value network using offline datasets to establish the LFR. Subsequently, it learns a reasonable adversarial policy for CBV based on a feasibility-dependent objective to create safety-critical scenarios with reasonable adversariality. Experiment results illustrate that the scenarios generated by FREA contain considerable near-miss events while effectively reducing unavoidable collision events. Moreover, guided by the AV’s feasibility and the goal-based adversarial reward, FREA demonstrates enhanced robustness and generalization ability in AV testing across various environments and surrogate AV methods.
Limitation. With all the advantages of FREA discussed, some limitations of our method also suggest directions for future work. As detailed in Section C.1, the LFR of AV exhibits instability under different hyperparameters. This instability is primarily due to the insufficient size of the offline dataset, a concern also highlighted in [15], which adversely affects the accuracy of LFR assessments. In addition, FREA faces challenges with traffic rule violations, as discussed in Section F.2. These violations are expected due to the absence of map information in the agent’s state. Integrating traffic rules into the policy training of the CBV is a crucial future direction for our research.
@@ A.1 Feasibility-depend Advantage Function
In [13], the feasibility-dependent advantage function relates to $A_{r}^{\pi_{\theta_{k}}}(s,a)$ and $A_{h}^{\pi_{\theta_{k}}}(s,a)$. Since the LFR of the AV is pre-calculated and independent of the policy $\pi$, we can substitute $A_{h}^{\pi_{\theta_{k}}}(s,a)$ with $A_{h}^{*}(s,a)$. Thus, the combined advantage function is defined as:
$\overline{A^{\pi_{\theta_{k}}}}(s,a)=A_{r}^{\pi_{\theta_{k}}}(s,a)+\lambda_{\xi}(s)\cdot A_{h}^{*}(s,a),$ (10)
@@
$\overline{A^{\pi_{\theta_{k}}}}(s,a)=A_{r}^{\pi_{\theta_{k}}}(s,a)\cdot\mathbb{I}_{s\in\mathcal{S}_{f}^{*}}+A_{h}^{*}(s,a)\cdot\mathbb{I}_{s\notin\mathcal{S}_{f}^{*}}$ (11)
@@
$\overline{A^{\pi_{\theta_{k}}}}(s,a)=A_{r}^{\pi_{\theta_{k}}}(s,a)\cdot\mathbb{I}_{V^{\ast}_{h}\left(g\left(s\right)\right)\leq 0}+A_{h}^{*}\left(s,a\right)\cdot\mathbb{I}_{V^{\ast}_{h}\left(g\left(s\right)\right)>0}$ (12)
@@ A.2 Proof of Lemma 1
According to Eq. 2, $Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}})$ is solely dependent on $h(\cdot)$, which in turn is influenced by $s_{t}^{\text{AV}},t\in\mathbb{N}$. The action $a^{\text{AV}}_{t}$ influences only the subsequent state $s^{\text{AV}}_{t+1}$ in conjunction with the environment transition function. Therefore, in a deterministic environment, this relationship can be described as follows:
$Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}}):=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{0}^{\text{AV}}\right),h\left(s_{t+1}^{\text{AV}}\right)\right\},\\ s_{0}^{\text{AV}}=s^{\text{AV}},a_{0}^{\text{AV}}=a^{\text{AV}},a_{t+1}^{\text{AV}}\sim{\pi^{\text{AV}}}\left(\cdot\mid s_{t+1}^{\text{AV}}\right),s_{t+1}^{\text{AV}}=P\left(s_{t}^{\text{AV}},a_{t}^{\text{AV}}\right),$ (13)
where $P\left(\cdot\right)$ denotes the deterministic transition dynamics. However, during the training of our CBV method, the policy varies significantly, impacting $s_{t+1}^{\text{AV}}$ through both the AV’s action $a_{t}^{\text{AV}}$ and the CBV’s action $a_{t}$. Given that other BVs adhere to a consistent rule-based policy, we can reasonably assume that the environment, excluding the CBV and AV, is deterministic. Thus, we can redefine the equation as follows:
$Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}},s,a):=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{0}^{\text{AV}}\right),h\left(s_{t+1}^{\text{AV}}\right)\right\},\\ s_{0}^{\text{AV}}=s^{\text{AV}},a_{0}^{\text{AV}}=a^{\text{AV}},s_{0}=s,a_{0}=a,s_{t+1}^{\text{AV}}=P\left(s_{t}^{\text{AV}},a_{t}^{\text{AV}},s_{t},a_{t}\right),s_{t+1}=g\left(s_{t+1}^{\text{AV}}\right)\\ a_{t+1}^{\text{AV}}\sim{\pi^{\text{AV}}}\left(\cdot\mid s_{t+1}^{\text{AV}}\right),a_{t+1}\sim{\pi^{\text{CBV}}\left(\cdot\mid s_{t+1}\right)},$ (14)
@@ 1. Case1:
    if $h\left({s^{\text{AV}}}^{\prime}\right)\geq h\left(s^{\text{AV}}\right)$, then:
    $Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}},s,a)=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{0}^{\text{AV}}\right),h\left(s_{t+1}^{\text{AV}}\right)\right\}=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{t+1}^{\text{AV}}\right)\right\}=V_{h}^{*}\left({s^{\text{AV}}}^{\prime}\right)$ (15)
@@ 2. Case2:
    if $h\left({s^{\text{AV}}}^{\prime}\right)<h\left(s^{\text{AV}}\right)$, then:
    $Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}},s,a)=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{0}^{\text{AV}}\right),h\left(s_{t+1}^{\text{AV}}\right)\right\}=\max\{h\left(s^{\text{AV}}\right),V^{*}_{h}\left({s^{\text{AV}}}^{\prime}\right)\}$ (16)
Building upon Eqs. 15 and 16, we conclude in Eq. 17 that $Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}},s,a)$ primarily depends on the states of the AV. The actions of both the AV and the CBV mainly influence state transitions.
$Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}},s,a)=Q_{h}^{*}(s^{\text{AV}},{s^{\text{AV}}}^{\prime})=\begin{cases}V^{*}_{h}\left({s^{\text{AV}}}^{\prime}\right)&h\left({s^{\text{AV}}}^{\prime}\right)\geq h\left(s^{\text{AV}}\right)\\ \max\{h\left(s^{\text{AV}}\right),V^{*}_{h}\left({s^{\text{AV}}}^{\prime}\right)\}&h\left({s^{\text{AV}}}^{\prime}\right)<h\left(s^{\text{AV}}\right)\end{cases}$ (17)
Thus, through the viewpoint transformation function $g(\cdot)$ and a deterministic AV policy $\pi^{\text{AV}}$, we derive the optimal feasible action-value function, taking into account both the state and action of the CBV:
$Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}})$ $=Q_{h}^{*}\left(s^{\text{AV}},a^{\text{AV}},s,a\right)$
$=Q_{h}^{*}(s^{\text{AV}},{s^{\text{AV}}}^{\prime})$
$=Q_{h}^{*}\left(g\left(s\right),g\left(s^{\prime}\right)\right)$
=$=\begin{cases}V^{*}_{h}\left(g\left(s^{\prime}\right)\right)&h\left(g\left(s^{\prime}\right)\right)\geq h\left(g\left(s\right)\right)\\ \max\{h\left(g\left(s\right)\right),V^{*}_{h}\left(g\left(s^{\prime}\right)\right)\}&h\left(g\left(s^{\prime}\right)\right)<h\left(g\left(s\right)\right)\end{cases}$ (18)

### v2

@@ 4 Experiments
Following the experimental setup (Section 4.1), we conducted several experiments to affirm: The Largest Feasible Region (LFR) of the AV is reasonable through offline training (Section 4.2); FREA effectively generates near-miss events while maintaining AV’s feasibility (Section 4.3); FREA shows generalization in AV testing across various surrogate AVs and environments (Section 4.4); and FREA continuously creates adversarial yet AV-feasible scenarios in complex traffic (Section 4.5).
@@ 4.1 Experiment Setup
Environment. We developed our environment using the Carla Simulator [24] in accordance with SafeBench [25], and combined “Scenario6” and “Scenario7” into “Scenario9”, spanning across “Town05” and “Town02”. In “Scenario9”, the AV navigates predefined routes through various junctions, avoiding randomly placed background vehicles (BVs). To implement reasonable adversarial attacks, we applied the FREA method to Critical Background Vehicles (CBVs), attacking the surrogate AV with a goal-based adversarial reward to create adversarial yet AV-feasible, safety-critical scenarios. More experimental details are provided in Appendix B.
@@ 4.2 Largest Feasible Region Learning of AV
According to the loss functions for the feasible value functions detailed in Eqs. 3 and 4, learning incorporates the constraint violation function $h(s^{\textnormal{AV}})$. A straightforward definition of $h(s^{\textnormal{AV}})$ is $\min_{i\in\{1,\ldots,N\}}d(\textnormal{AV},\textnormal{BV}^{i})$, where $d(\textnormal{AV},\textnormal{BV}^{i})$ represents the minimum distance between the bounding boxes of AV and ${\textnormal{BV}}^{i}$. However, since the minimum distance between boxes is non-negative, we adopt a sparse design for $h(s^{\textnormal{AV}})$, which helps to maintain a stable training balance of positive and negative samples, as advised by [15]:
$h(s^{\textnormal{AV}}):=\begin{cases}-1&\text{if }\min_{i\in\{1,\ldots,N\}}d(\textnormal{AV},\textnormal{BV}^{i})>d_{th}\\ M&\text{if }\min_{i\in\{1,\ldots,N\}}d(\textnormal{AV},\textnormal{BV}^{i})\leq d_{th}\end{cases},$ (10)
@@
Figure 2: Visualization of well-trained LFR $\left\{s^{\textnormal{AV}}|V^{\ast}_{h}(s^{\textnormal{AV}})\leq 0\right\}$ in various traffic scenarios.
Figure 2 depicts the well-trained LFR of the AV in varying traffic scenarios. Figure 2(a) and (b) reveal that the infeasible region expands as the AV’s speed increases, indicating higher collision risks. Figure 2(c) and (d) confirm the LFR’s consistency across scenarios. These cases show the pre-trained LFR’s reliability under various scenarios, providing a solid basis for CBV policy training.
@@ 4.3 Quantitative Analysis of Generated Scenarios
Evaluation of Near-miss Events. Previous research indicates that all safety-related incidents share common origins with near-misses [27, 28, 29], highlighting the importance of analyzing near-miss events in the generation of safety-critical scenarios. Consistent with [4], we use time-to-collision (TTC) in [30] and post-encroachment time (PET) in [4] as metrics to quantify these near-miss events.
Figure 3(a) and (b) show the distributions of TTC and PET across different CBV methods. The PPO produces the most adversarial scenarios across all maps. Conversely, FPPO-RS and FREA exhibit similar near-miss results but with a higher frequency of near-miss than Standard traffic flow. Given our focus on AV-feasible near-miss events, further analysis of AV’s feasibility is required.
Evaluation of AV’s Feasibility. While near-miss events are crucial for generating safety-critical scenarios, it is also vital to avoid over-collision events caused by excessive adversarial CBV policy. To ensure that FREA effectively minimizes over-collision events, we analyze the scenarios under different CBV methods with four metrics: (1) Collision Rate (CR); (2) Infeasible Ratio (IR); (3) Infeasible Distance (ID); and (4) Collision Severity, including collision speed and impulse. See Section E.2 for metric details and Section B.3 for detailed reimplementation of KING [5].
@@
As shown in Table 1 and Figure 4, without the feasibility constraint, the PPO and KING [5] lead to many severe collisions, indicative of excessive adversarial behavior. In contrast, under the feasibility constraint, both FPPO-RS and FREA significantly reduce the severity of these events. However, while addressing AV’s feasibility, FPPO-RS struggles with hard constraints, which leads to many infeasible events. Conversely, FREA adeptly balances adversarial intensity with AV’s feasibility, thereby achieving the least severe collisions.
@@
To evaluate the generalization of FREA across various surrogate AVs and environments in AV testing, we assessed the driving performance of PlanT [31] and Expert [26] under different CBVs that were pre-trained with specific surrogate AVs on “Town05” and “Town02”. The performance metrics follow the setting in [25], where the OS represents the overall score (details in Section E.3).
Table 2 illustrates that the PPO method, pre-trained with various surrogate AV methods, exhibits notable variance in AV testing. This variance suggests that, without feasibility constraints, the CBV’s adversarial policy closely depends on the specific surrogate AV method. In contrast, FREA demonstrates remarkable stability in AV testing. This stability is primarily due to goal-based adversarial rewards and feasibility constraints, both independent of the specific surrogate AV method employed. Furthermore, FREA consistently aligns its relative AV testing results with standard traffic flows, suggesting that FREA facilitates safety-critical scenario generation while ensuring unbiased AV testing.
@@ 5 Conclusion
In conclusion, FREA employs the concept of the Largest Feasible Region (LFR) of the AV to establish the upper bound of adversariality in safety-critical scenario generation. This approach aims to maximize adversariality while ensuring AV’s feasibility. Specifically, FREA employs a two-stage framework that begins by pre-training the AV’s optimal feasible value network using offline datasets to establish the LFR. Subsequently, it learns a reasonable adversarial policy for Critical Background Vehicles (CBVs) based on a feasibility-dependent adversarial objective, thereby creating adversarial yet AV-feasible safety-critical scenarios. Experimental results illustrate that the scenarios generated by FREA contain considerable near-miss events while effectively reducing unavoidable collision events. Moreover, guided by the AV’s feasibility and the goal-based adversarial reward, FREA demonstrates enhanced robustness and generalization ability in AV testing across various environments and surrogate AV methods.
Limitation. With all the advantages of FREA discussed, some limitations of our method also suggest directions for future work. As detailed in Section C.1, the LFR of AV exhibits instability under different hyperparameters. This instability is primarily due to the limited offline data size, a concern also highlighted in [15], which adversely affects the accuracy of LFR assessments. In addition, FREA lacks certain traffic common sense and is prone to violating traffic regulations, as discussed in Section F.2. In the future, integrating intelligent agents like LLMs to address these deficiencies and ensure compliance with traffic rules will be an important research direction.
Acknowledgments
This research was supported by the National Natural Science Foundation of China under Grant 52221005 and the Tsinghua-Toyota Joint Research Institute Interdisciplinary Project under Grant 20223930039.
@@ A.1 Feasibility-depend Advantage Function
In [13], the feasibility-dependent advantage function relates to $A_{r}^{\pi_{\theta_{k}}}(s,a)$ and $A_{h}^{\pi_{\theta_{k}}}(s,a)$. However, as we use the Largest Feasible Region (LFR) of the AV as the guidance to train the CBV policy and the predetermined LFR is independent of the CBV policy $\pi$, $A_{h}^{\pi_{\theta_{k}}}(s,a)$ can be substituted with $A_{h}^{*}(s^{\text{AV}},a^{\text{AV}})$. Thus, the combined advantage function is defined as:
$A^{\pi_{\theta_{k}}}(s,a)=A_{r}^{\pi_{\theta_{k}}}(s,a)+\lambda_{\xi}(s)\cdot A_{h}^{*}(s^{\text{AV}},a^{\text{AV}}),$ (11)
@@
$A^{\pi_{\theta_{k}}}(s,a)=A_{r}^{\pi_{\theta_{k}}}(s,a)\cdot\mathbb{I}_{s\in\mathcal{S}_{f}^{*}}+A_{h}^{*}(s^{\text{AV}},a^{\text{AV}})\cdot\mathbb{I}_{s\notin\mathcal{S}_{f}^{*}}$ (12)
@@
$A^{\pi_{\theta_{k}}}(s,a)=A_{r}^{\pi_{\theta_{k}}}(s,a)\cdot\mathbb{I}_{V^{\ast}_{h}\left(g\left(s\right)\right)\leq 0}+A_{h}^{*}(s^{\text{AV}},a^{\text{AV}})\cdot\mathbb{I}_{V^{\ast}_{h}\left(g\left(s\right)\right)>0}$ (13)
@@ A.2 Proof of Lemma 1
Given that other BVs adhere to a consistent rule-based policy, we can reasonably assume that the environment, excluding the CBV and AV, is deterministic. Thus, we can redefine Eq. 2 as follows:
$Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}}):=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{0}^{\text{AV}}\right),h\left(s_{t+1}^{\text{AV}}\right)\right\},\\ s_{0}^{\text{AV}}=s^{\text{AV}},a_{0}^{\text{AV}}=a^{\text{AV}},s_{0}=s,a_{0}=a,s_{t+1}^{\text{AV}}=P\left(s_{t}^{\text{AV}},a_{t}^{\text{AV}},s_{t},a_{t}\right),s_{t+1}=g\left(s_{t+1}^{\text{AV}}\right)\\ a_{t+1}^{\text{AV}}\sim{\pi^{\text{AV}}}\left(\cdot\mid s_{t+1}^{\text{AV}}\right),a_{t+1}\sim{\pi^{\text{CBV}}\left(\cdot\mid s_{t+1}\right)},$ (14)
where $s$ and $a$ are the state-action pair of the CBV and $P$ is the deterministic transition function excluding the CBV and AV.
@@ 1. Case1:
    if $h({s^{\text{AV}}}^{\prime})\geq h(s^{\text{AV}})$, then:
    $Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}})=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{0}^{\text{AV}}\right),h\left(s_{t+1}^{\text{AV}}\right)\right\}=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{t+1}^{\text{AV}}\right)\right\}=V_{h}^{*}({s^{\text{AV}}}^{\prime})$ (15)
@@ 2. Case2:
    if $h({s^{\text{AV}}}^{\prime})<h(s^{\text{AV}})$, then:
    $Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}})=\min_{\pi^{\text{AV}}}\max_{t\in\mathbb{N}}\left\{h\left(s_{0}^{\text{AV}}\right),h\left(s_{t+1}^{\text{AV}}\right)\right\}=\max\{h(s^{\text{AV}}),V^{*}_{h}({s^{\text{AV}}}^{\prime})\}$ (16)
Building upon Eqs. 15 and 16, we conclude in Eq. 17 that $Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}})$ primarily depends on the states of the AV. The actions of both the AV and the CBV mainly influence state transitions.
$Q_{h}^{*}(s^{\text{AV}},a^{\text{AV}})=\begin{cases}V^{*}_{h}({s^{\text{AV}}}^{\prime})&h({s^{\text{AV}}}^{\prime})\geq h(s^{\text{AV}})\\ \max\{h(s^{\text{AV}}),V^{*}_{h}({s^{\text{AV}}}^{\prime})\}&h({s^{\text{AV}}}^{\prime})<h(s^{\text{AV}})\end{cases}$ (17)